OpenStudy (anonymous):

solve the differential equation using the variation of parameters method \[y''+y=sec^2x,0\le x\le \frac{\pi}2\] \[r^2+1=0\]

OpenStudy (anonymous):

did you get that problem from the other day sorted out?

OpenStudy (anonymous):

y" +y =t*e^t iirc?

OpenStudy (anonymous):

at first I had r^2+r=0 and that gave me the integral of \[\int\frac{e^{-x}sec^2x}{e^{-x}}dx+e^{-x}\int \frac{sec^2x}{e^{-x}}dx\]

OpenStudy (anonymous):

but that is apparently wrong

OpenStudy (anonymous):

iirc? what does that mean?

OpenStudy (anonymous):

if I recall correctly

OpenStudy (anonymous):

\[r_1=r_2=\pm1i\]

OpenStudy (anonymous):

\[y_c=c_1cosx+c_2sinx\]

OpenStudy (anonymous):

\[W(y_1,y_2)=cos^2x+sin^2x\]

OpenStudy (anonymous):

1

OpenStudy (anonymous):

nice! \[y_p=-cosx\int sinxsec^2xdx+sinx \int cosxsec^2xdx\]

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[y=y_c+y_p=c_1cosx+c_2sinx-cosxsecx+\int cosxsec^2xdx\]

OpenStudy (anonymous):

missing a 'sin' on that last term...

OpenStudy (anonymous):

@mathsofiya , check integration seems incomplete

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

@Algebraic! , didnt really see the reason for "lol", there are couple of things missing here

OpenStudy (anonymous):

"integration seems incomplete"

OpenStudy (anonymous):

what tipped you off?

OpenStudy (anonymous):

here I thought you were making a funny and you were somehow serious...

OpenStudy (anonymous):

a simple look suggests that last term carries an Integration sign whereas the second last term does not, it cant be both , either we have an integral sign or we dont. Anyways its better to teach something correctly : )

OpenStudy (anonymous):

where's a sine missing?

OpenStudy (anonymous):

sinx∫cosxsec^2x dx -> ∫cosxsec^2x dx

OpenStudy (anonymous):

very true \[y=y_c+y_p=c_1cosx+c_2sinx-cosxsecx+sinx\int cosxsec^2xdx\]

OpenStudy (anonymous):

is this really the solution to the integral? http://www.wolframalpha.com/input/?i=integral+of+cosxsec^2x

OpenStudy (anonymous):

most people use ln(tan(x) +sec(x)) for brevity

OpenStudy (anonymous):

programs always give the other form of the answer; not sure why, probably to confound students.